With what minimum speed a particle be projected from the origin so that it is able to pass through a given point (30,40)

Grade:12th pass

1 Answers

Khimraj
5 years ago
For min velocity point (30,40) will be max height point.
Then tan\Theta = 40/30 = 4/3
H = 40 = u2sin2\Theta/2g
40 = u2(4/5)2/2*10
u2 = 800*25/16 = 1250
then u = 25\sqrt{2} m/s
Hope it clears.
If you like answerthen please approve it.